Given that the roots of a quadratic equation ax^2 + bx+c = 0 are   m and n ,      Write down the 2 roots ,   m  and n  , in terms of a, b, c.

Grade:

1 Answers

Nikhil Cherukupalli
12 years ago

 We know that, m+n = -b/a and m.n = c/a ..

Now,(m-n)2 = (m+n)2 - 4mn => m-n = √(m+n)2 - 4mn

Adding (m+n) and (m-n),we get m= [-b/a + √(m+n)2 - 4mn]/2

Similarly,subtract (m-n) from (m+n) to get n = [-b/a - √(m+n)2 -4mn]/2